### v2

# A note on real algebraic maps which are topologically special generic maps

Abstract.
Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps. A special generic map is, roughly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
This paper constructs real algebraic maps topologically special generic maps whose images are smoothly embedded manifolds.
We are also interested in construction of explicit and meaningful smooth maps in differential topology and recently ones in real algebraic geometry. This has been an important and difficult problem.
In such stories, we have previously constructed real algebraic maps topologically regarded as special generic maps. This paper is a kind of additional short remark on such maps.
Key words and phrases:
(Non-singular) real algebraic manifolds and real algebraic maps. Smooth maps. Special generic maps.
2020 Mathematics Subject Classification: Primary 14P05, 14P25, 57R45, 58C05. Secondary 57R19.
1. Introduction.
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [4, 7, 28] are related pioneering studies. For Morse functions, see [26] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has also another nice history. [24, 35] show classical and important studies and [1, 2, 16, 17] explain fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
Celebrated theory by Nash, followed by Tognoli, says that smooth closed manifolds are non-singular real algebraic manifolds. They are also shown to be the zero sets of some real polynomial maps. We can also approximate smooth maps between non-singular real algebraic manifolds by real algebraic ones. It is also important and difficult to have real algebraic functions and maps explicitly.
For example, [19, 25, 33] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [10], followed by [14] by the author himself. We present results of a similar type which are also new.
1.1. Smooth manifolds and maps.
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [23] for example. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds. It is also the Riemannian manifold with the standard Euclidean metric. Let $\mathbb{R}:={\mathbb{R}}^{1}$. For $x\in{\mathbb{R}}^{k}$, $||x||\geq 0$ is defined as the canonically defined distance between $x$ and the origin 00. This is also regarded as a (non-singular) real algebraic manifold or the $k$-dimensional real affine space. $S^{k}:=\{x\in{\mathbb{R}}^{k+1}\mid||x||=1\}$ denotes the $k$-dimensional unit sphere, which is a $k$-dimensional smooth compact submanifold of ${\mathbb{R}}^{k+1}$ and has no boundary. It is connected for any positive integer $k\geq 1$. It is a discrete set with exactly two points for $k=0$. It is the zero set of the real polynomial $||x||^{2}={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface). $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk. It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$. Let $S(c)$ denote the set of all singular points of $c$. In our paper, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
Here, a real algebraic manifold means some union of connected components of the intersection of finitely many sets represented as the zero sets of some real polynomial maps. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. Note that we may say Nash manifolds and maps for example for such maps. However, we do not use such terminologies. [30] show related fundamental theory, arguments and some advanced theory.
1.2. Our main results.
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [8] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result.
Proposition 1 ([28]).
For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
  1. (1)
    We have an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ in such a way that $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
  2. (2)
    We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-N(\partial W_{f})$ whose fiber is the unit sphere.
Main Theorem 1.
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps on two suitably obtained mutually homeomorphic $m$-dimensional closed and connected manifolds $M$ and $M_{0}$ with the following properties.
  1. (1)
    $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
  2. (2)
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}$.
  3. (3)
    There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
The next section proves Main Theorems. The third section explains about meanings of our main result and related topics by some remarks.
Conflict of interest.
This work was also supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”. The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/). This is closely related to our study. Our study thanks this project for supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He is not employed there. However this helps our studies. Our study also thanks this.
Data availability.
Data essentially supporting our present study are all here.
2. On Main Theorems.
We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00. This is a main theorem of [28].
In [15], followed by [13], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example. Hereafter, for a manifold $X$, ${\rm Int}\ X$ denote the interior $X-\partial X$ of it.
Proposition 2 ([28]).
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do.
  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have this enjoying the following properties.
  1. (1)
    The restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ to $\bar{N}-N(\partial\bar{N})\subset\bar{N}-{\rm Int}\ N(\partial\bar{N})$ and the bundle over $\bar{N}-N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
  2. (2)
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ are isomorphic and naturally identified.
  3. (3)
    We also consider the restriction of bundle ${\bar{M}}_{\rm I}$ in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
Definition 1.
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[15] has first introduced such a class of special generic maps. We first name the class in this way first in [12]. Of course in our present paper, we do not need to understand this notion or [12, 15] well and we do not assume their arguments or results.
However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This gives another simplest example.
We add the following as additional main results.
Main Theorem 2.
We can have a pair of maps in Main Theorem 1 as follows.
  1. (1)
    $f$ is product-organized.
  2. (2)
    $M_{0}$ is also represented as the zero set of some real polynomial.
Main Theorem 3.
In Main Theorem 2, we can have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
  1. (1)
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” and have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and whose structure group is reduced to a group consisting of linear transformations of $D^{m-n+1}$: in fact a manifold homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined except for $m-n+1=k\neq 4$.
  2. (2)
    Under the structure group, the bundle of (1) here is in fact trivial.
  3. (3)
    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ diffeomorphic to a manifold obtained in the following steps.
    1. (a)
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ homeomorphic to $D^{m-n+2}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
    2. (b)
      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
    3. (c)
      Eliminate the corner.
    In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered.
A proof of Main Theorems.
We prove our main result. First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
This is a new ingredient in our related studies. Our arguments here thanks to [16]. Especially, ”Special Case 5” and ”Definition 14”.
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ by considering a suitable isotopy.
In addition, we can also get the embedding and the subset in such a way that the following are satisfied.
  • ${\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}\geq 0$.
  • ${\rm Int}\ {\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}>0$.
  • At each point of ${\mathbb{R}}^{n}-{\bar{N}}_{0}$, $f_{{\partial}_{j}}<0$ holds for exactly one $1\leq j\leq l$ and $f_{{\partial}_{j}}>0$ holds for the remaining polynomials.
STEP 2 Arguments from [10, 14] to have our smooth real algebraic map $f_{0}$.
We apply arguments from main ingredients of [10, 14].
We consider the zero set $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
In STEP 1, we have seen that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function. Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way. By considering the partial derivative for some variable $x_{i}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We apply the implicit function theorem to see that
$M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
STEP 3 Constructing a special generic map $f$.
For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
We construct a homeomorphism from $M_{0}$ onto another smooth manifold $M$. First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$ where $0\in{\mathbb{R}}^{m-n+1}$ here is the origin of course.
We can consider a smooth function $g:(0,\infty]\rightarrow(0,1]$ enjoying the following.
  • $g(t_{1})\leq g(t_{2})$ for $t_{1}<t_{2}$.
  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$
  • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the standard way.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
We have a homeomorphism onto another manifold $M$. This is smooth and homeomorphic to $M_{0}$. We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
The resulting map is also special generic by the construction and product organized. By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2. This completes the proof of Main Theorems 1 and 2.
We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
We first show Main Theorem 3 (1). First the bundle we discuss here is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1. More precisely, Proposition 2.1 of [28] is a corresponding result. It also refers to [22] there. The structure group is regarded as a group consisting of diffeomorphisms preserving the distance between each point and the origin $0\in{\Delta}^{m-n+1}$ where we can define the origin 00 for ${\Delta}^{m-n+1}$ not diffeomorphic to $D^{m-n+1}$. We explain about this. For an element ${\phi}_{{\Delta}^{m-n+1}}$ of such a group, we consider $\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)$ for $x\in{\Delta}^{m-n+1}$ and $0<t\leq 1$ where $x$ and $tx$ are canonically regarded as vectors. Here $||\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)||=||x||$. We consider the limit at $t=0$. For a fixed $x$, we have the value of the differential of ${\phi}_{{\Delta}^{m-n+1}}$ at $x$, seen naturally as a tangent vector at the origin of ${\Delta}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the structure group of a linear bundle whose fiber is $D^{m-n+1}$. This completes the proof of Main Theorem 3 (1).
We see that this bundle is trivial for the structure group.
$W_{0}$ is defined as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
and this can be also defined as
$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$.
We use the function $g$ before. We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule where we consider some elements as vectors.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$. By the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$, we can see that the previous bundle is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
By our arguments, we can also check Main Theorem 3 (3) easily. A result close to this is also presented in Theorem 1 (3) of [14]. ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there: only exposition on ”homotopy” is presented there.
This completes the proof. ∎
3. Meanings of our main result and remarks.
The first remark compares our main result to previous results by the author.
Remark 1.
In [10, 11, 14], we consider cases where polynomials are given beforehand. In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark and this does not affect main ingredients of [10, 11, 14]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
The next remark is related to structures of our new real algebraic maps.
Remark 2.
It seems that our new real algebraic maps are not special generic in general in the smooth category. It is important to know whether our maps are special generic or not. It is also difficult.
The following is on another problem.
Remark 3.
As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing canonical projections. Some are closely related to Remark 1.
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. [26] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006 [29] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [18, 20] show explicit answers. [9] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[10] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [3, 31, 32] are also closely related, help the author to do such studies, and motivate us to study further.
The following is a remark on Main Theorem 3.
Remark 4.
As another method, we can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
4. Acknowledgement.
This new study is based on studies on special generic maps and explicit construction of real algebraic functions and maps of the author. One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry.
The author would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference. He knows singularity theory and applications to geometry widely. Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [12] of the author. He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [34, 36], studies [5, 6] on existence, [27], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [8] for related fundamental and advanced theory from singularity theory for example).
The author would like to thank Atsufumi Honda for asking questions and related informal discussions at Kyushu University after the conference on this theorem, which was an event independent of the previous discussion and held on 2023/12/8. It is also on non-existence of special generic maps on projective spaces and interesting problems from the viewpoint of differential geometry. As a researcher on singularity theory and applications to differential geometry, he says that Reeb’s theorem on Morse functions and special generic maps seem to be closely related to minimizing the total curvatures of curves, surfaces and Riemannian manifolds.
Last, in the informal discussions with Takahiro Yamamoto, the author also says that he is now studying differential geometry and real algebraic geometry further to understand manifolds in related categories, which are finer than the smooth category. We have also discussed the discussions with Atsufumi Honda. He has also commented that special generic maps obtain some nice global structures of manifolds.
In conclusion, we have enjoyed the discussions to present this study.