### v5

# Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres

Abstract.
Morse functions with exactly two singular points on spheres and canonical projections of spheres are generalized to special generic maps. A special generic map is, roughly, a smooth map represented as the composition of a surjection onto a manifold with an immersion of codimension 00. Preimages of this surjections are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary.
We construct real algebraic maps topologically seen as special generic maps whose images are smoothly embedded manifolds. Canonical projections of unit spheres are generalized here. We are interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry. Obtaining them explicitly is important and difficult. In the stories, we previously constructed some real algebraic maps topologically special generic maps. We construct such maps in more general situations and give additional remarks.
Key words and phrases:
(Non-singular) real algebraic manifolds and real algebraic maps. Real algebraic hypersurfaces. Hilbert’s 16th problem. Smooth maps. Special generic maps.
2020 Mathematics Subject Classification: Primary 14P05, 14P25, 57R45, 58C05. Secondary 57R19.
1. Introduction.
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [5, 9, 44] are related pioneering studies. For Morse functions, see [41] for a pioneering study and [36] for systematic exposition on fundamental notions and arguments. For higher dimensional versions of theory of Morse functions, [52, 55] are pionnering studies.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has another nice history. [39, 53] show classical and important studies and [2, 3, 21, 22] present fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of) spheres.
Celebrated theory by Nash, followed by Tognoli, says that smooth closed manifolds are non-singular real algebraic manifolds. They are also shown to be the zero sets of some real polynomial maps. We can also approximate smooth maps between non-singular real algebraic manifolds by real algebraic ones. It is also important and difficult to have real algebraic functions and maps explicitly.
For example, [25, 40, 51] show explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We consider another generalization for projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [15], followed by [18] by the author himself. We present new related result. More precisely, we construct new explicit maps on some real algebraic hypersurfaces. As one of related future work, [10, 54] present closely related long-standing studies on construction of real algebraic curves in the plane and general real algebraic hypersurfaces (Remark 5).
1.1. Smooth manifolds and maps.
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [38] for example. Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ the interior of it.
Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds. It is also the Riemannian manifold with the standard Euclidean metric. Let $\mathbb{R}:={\mathbb{R}}^{1}$. For $x\in{\mathbb{R}}^{k}$, $||x||\geq 0$ is defined as the canonically defined distance between $x$ and the origin 00. This is also regarded as a (non-singular) real algebraic manifold or the $k$-dimensional real affine space. $S^{k}:=\{x\in{\mathbb{R}}^{k+1}\mid||x||=1\}$ denotes the $k$-dimensional unit sphere, which is a $k$-dimensional smooth compact submanifold of ${\mathbb{R}}^{k+1}$ and has no boundary. It is connected for any positive integer $k\geq 1$. It is a discrete set with exactly two points for $k=0$. It is the zero set of the real polynomial $||x||^{2}-1={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}-1$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface). $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk. It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
Let $c:X\rightarrow Y$ be a differentiable map. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$. Let $S(c)$ denote the set of all singular points of $c$. Here, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. The canonical projection of the unit sphere $S^{k-1}$ is its restriction.
Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. We may say Nash manifolds and maps for example for such maps. [46] shows related fundamental theory, arguments and some advanced theory.
1.2. Our main result and the structure of our paper.
Spaces of smooth maps between two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies for each non-negative integer $r$. They respect differentials of degrees at most $r$. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [11] for systematic exposition. Diffeomorphisms are smooth maps with no singular points (of them) and being homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
We present a fundamental theorem on structures of special generic maps as Proposition 1. Using the notation, we present our new real algebraic maps as Main Theorem 1. This is proven in the next section with our additional result Main Theorems 2 and 3. Main Theorem 2 is a kind of remarks on Main Theorem 1. Main Theorem 3 is a more generalized result. The third section is on meanings of our achievement and related topics.
Proposition 1 ([44]).
Let $m>n\geq 1$ be integers. For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
  1. (1)
    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ and $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
  2. (2)
    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
Main Theorem 1.
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then there exist a pair $(M,M_{0})$ of two mutually diffeomorphic $m$-dimensional closed and connected manifolds and a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps there and the following properties are enjoyed.
  1. (1)
    The map $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$. The map $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ being also the zero set of a real polynomial.
  2. (2)
    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f_{0}(S_{0}(f_{0})))=f(S(f))$.
This paper is regarded as an improved version of [19] and based on our new discovery and idea.
Conflict of interest.
This work has been supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”. The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/). This is closely related to our study. Our study thanks this project for supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He is not employed there. However this helps our studies. Our study also thanks this.
Data availability.
Data essentially supporting our present study are all here.
2. On Main Theorems.
For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, let the closed interval denoted by $[a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}\leq x\leq a_{2}\}$ as is the standard notation. Let the open interval denoted by $(a_{1},a_{2}):={\rm Int}\ [a_{1},a_{2}]$ as is the standard case. Similarly, we also define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$, for example and for $a\in\mathbb{R}$, let us define $[a,\infty):=\{x\in\mathbb{R}\mid x\geq a\}$ for example.
We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00. This is a main theorem of [44]. In [20], followed by [17], the author has redefined several notions and consider specific reconstruction. We also abuse these notions for example.
Proposition 2 ([44]).
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do. We can also abuse this in the case $m=n$ and we do.
  • Consider a small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$ and this is the first object.
  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$ and this is the second object.
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $(\partial\bar{N})\times[0,1]$. Furthermore, each point $x\in\partial\bar{N}$ is identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$. In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}$ is identified with the space $(\partial\bar{N})\times\{1\}$ naturally and it is the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The third object is a bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
Then there exist an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have the manifold and the map enjoying the following properties.
  1. (1)
    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})=W_{f}-{\rm Int}\ N(\partial W_{f})$ in Proposition 1 are isomorphic and naturally identified.
  2. (2)
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified.
  3. (3)
    We also consider the restriction of the smooth bundle in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ for ${\bar{f}}_{N}$.
Note that in Proposition 1, we cannot set ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].
Definition 1.
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[20] has first introduced such a class of special generic maps. We first name the class in this way in [17]. Of course in our present paper, we do not need to understand this notion or [17, 20] well. We do not assume their arguments or results.
However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can reconstruct a product-organized special generic map into ${\mathbb{R}}^{n}$ there from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This gives another simplest example.
We add the following new result. This is also a kind of remarks on Main Theorem 1.
Main Theorem 2.
In Main Theorem 1, $f$ and $f_{0}$ can be obtained as product-organized special generic maps.
A proof of Main Theorems 1 and 2.
STEPs 1–3 are main ingredients.
STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
This is a new ingredient in our related studies. Our arguments here thanks to [21]. Especially, ”Special Case 5” and ”Discussion 14”.
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [23]. For this approximation, see also [4, 8].
This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that the manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
In addition, we can also get the embedding and the subset with the following properties.
  • ${\bar{N}}_{0}={\bigcap}_{j=1}^{l}\{x\in{\mathbb{R}}^{n}\mid f_{{\partial}_{j}}(x)\geq 0\}$.
  • ${\rm Int}\ {\bar{N}}_{0}={\bigcap}_{j=1}^{l}\{x\in{\mathbb{R}}^{n}\mid f_{{\partial}_{j}}(x)>0\}$.
  • At each point $x$ of ${\mathbb{R}}^{n}-{\bar{N}}_{0}$, the relation $f_{{\partial}_{j_{x}}}(x)<0$ holds for exactly one integer $1\leq j_{x}\leq l$ and the relation $f_{{\partial}_{j}}(x)>0$ holds for the remaining polynomials.
STEP 2 Arguments from [15, 19] to have our smooth real algebraic map $f_{0}$.
We apply arguments from main ingredients of [15, 19].
We consider the zero set $S_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
The last part of STEP 1 implies the fact that $S_{0}=\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$. We apply the implicit function theorem to see that this set is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a value which is not 00. We discuss this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ is not 00 for some variable $x_{j_{0}}$, since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary. Remember that this is also obtained via an isotopy in the $C^{1}$ class. For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$. The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is not 00.
We have checked that the set $S_{0}$ is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$. Put $M_{0}:=S_{0}$. Our desired map $f_{0}$ is defined as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
STEP 3 Is $f_{0}$ special generic?
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider our identification in the following way thanks to arguments related to ”Discussion 14” in [21].
  • The manifold $(\partial{\bar{N}}_{0})\times\{0\}$ is identified with $(\partial N(\partial{\bar{N}}_{0}))\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
  • The manifold $(\partial{\bar{N}}_{0})\times\{t_{0}\}\subset(\partial{\bar{N}}_{0})\times(0,1]$ is identified with the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}})-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$. We have ${\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=at_{0}={\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ at each point $(x,y)\in(((\partial{\bar{N}}_{0})\times\{t_{0}\})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}\subset({\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ ($y=(y_{1},\cdots,y_{m-n+1})$).
This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of a Morse function for a natural height of an ($m-n+1$)-dimensional disk and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function. Some $y_{j}$ is regarded as the variable of the target of the function here.
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ onto $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$. By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
These arguments imply that $f_{0}$ is a product-organized special generic map.
We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3.
By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
This completes the proof.
We show Main Theorem 3. This is a kind of generalizations of Main Theorem 2.
Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly. See [35] and see also [13, 47, 48] for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
Main Theorem 3.
Suppose that a situation same as one of Main Theorem 1 is given.
Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following. We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $y$.
  1. (1)
    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
  2. (2)
    $F_{0,S^{m-n}}(0)=0$. $F_{0,S^{m-n}}(t)>0$ for $t>0$ and $F_{0,S^{m-n}}(t)<0$ for $t<0$. The value of the derivative of $F_{0,S^{m-n}}$ at 00 is not zero.
  3. (3)
    ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$. $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)\geq 0$ for $y\in{\mathbb{R}}^{m-n+1}$.
  4. (4)
    We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively. The zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the polynomial is non-singular and it does not contain the origin $0\in{\mathbb{R}}^{m-n+1}$. It is also homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
Then we can have a statement of Main Theorem 1 by defining
$M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2.
Explicitly, we can also have the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
  1. (5)
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$. Furthermore, the smooth bundle is a trivial smooth bundle.
  2. (6)
    The non-singular real algebraic manifold $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps. Here ${\Delta}^{k}$ is for a $k$-dimensional smooth manifold homeomorphic to $D^{k}$ ($k\geq 1$). This is diffeomorphic to $D^{k}$ for $k\neq 4,5$. This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if the boundary is diffeomorphic to $S^{k-1}$.
    1. (a)
      Prepare the product $\bar{N}\times{\Delta}^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$.
    2. (b)
      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and ${\Delta}^{m-n+1}$: here ${\Delta}^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
    3. (c)
      Eliminate the corner.
Hereafter, for a real number $t$ and $x\in{\mathbb{R}}^{k}$, we use the notation respecting the structure of the canonical vector space for ${\mathbb{R}}^{k}$ such as $tx\in{\mathbb{R}^{k}}$. The vector $tx$ is also a point of ${\mathbb{R}}^{k}$ of course.
A proof of Main Theorem 3.
We can argue as STEP 1 in the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$. We choose another positive number $T_{0}<a$. We define $M_{0}$ as presented in the statement of Main Theorem 3.
We also define the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ as the composition of the canonical embedding to ${\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}$ as in Main Theorems 1 and 2 and their proof.
We present arguments similar to ones of STEP 2 in the proof of Main Theorems 1 and 2. Here we also respect the assumptions (1)–(4) here.
First we also have the relation
$M_{0}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
by our definitions similarly. We show that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. We discuss this more precisely. We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$. According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$. It is assumed in (2) that $F_{0,S^{m-n}}(t)>0$ for $t>0$. We can know that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 in the proof of Main Theorems 1 and 2. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the value in the case of Main Theorems 1 and 2. The value in the case of Main Theorems 1 and 2 means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the variable $x_{j_{0}}$. The resulting value is not 00.
We can argue similarly to see that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
We present an argument similar to one in STEP 3 in the proof of Main Theorems 1 and 2. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$. For example, we can consider this as the collar neighborhood of STEP 3 in the proof. We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$. We use the natural identification mapping $(x,0)$ to $(x,0)$ here. According to the property (3) for example, we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
We can consider a smooth function $g:(0,\infty)\rightarrow(0,1]$ enjoying the following.
  • $g(t_{1})\leq g(t_{2})$ for $t_{1}<t_{2}$.
  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$.
  • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
$(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$
according to the following rule.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
Thus we have a homeomorphism from the smooth manifold $M_{0}$ onto a smooth manifold $M$. The restriction of this to the set $M_{0}-(({\partial\bar{N_{0}}})\times\{0\})$ gives a diffeomorphism.
We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized. We define a special generic map $f$ as this map. An essentially same argument also appears in Main Theorems 1 and 2 and their proof.
We define a new set $W_{0}\subset{\mathbb{R}}^{m+1}$ as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
and by our assumption, this can be also defined as
$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$.
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
This rule is in fact same as the previous one. This gives a diffeomorphism. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a homeomorphism from $W_{0}$. The image of the homeomorphism is the disjoint union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
From this we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$. The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$. In other words, it is also regarded as a ”canonical projection of $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ onto ${\rm Int}\ {\bar{N}}_{0}$”.
We can define another set
$M_{0,t^{\prime}}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$
for $0<t^{\prime}\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. We present several properties based on our presented arguments. We can also know $M_{0,t^{\prime}}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$. For distinct real numbers $t^{\prime}=t_{1},t_{2}$ here, the sets $M_{0,t_{1}}-((\partial\bar{N_{0}})\times\{0\})$ and $M_{0,t_{2}}-((\partial\bar{N_{0}})\times\{0\})$ are disjoint. These sets $M_{0,t^{\prime}}-((\partial\bar{N_{0}})\times\{0\})$ are $m$-dimensional smooth manifolds with no boundaries. They are seen as the graphs of smooth functions as $M_{0}-((\partial\bar{N_{0}})\times\{0\})$ and diffeomorphic to $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. Here some variable $y_{j}$ is regarded as the variable of the target of the function. We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold. This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$.
From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}})-((\partial{\bar{N}}_{0})\times\{0\}))$.
We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). For this, first, remember that the projection is the restriction of the canonical projection ${\pi}_{m+1,n}$. Remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function. Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example. The relation $S(f_{0})=(\partial{\bar{N}}_{0})\times\{0\}$ also holds.
In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle. Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
Last we consider a small isotopy as in the proof of Main Theorems 1 and 2. This completes the proof of the statement of Main Theorem 1.
We can also check the property (5) by investigating arguments of STEP 3 of our proof of Main Theorems 1 and 2 and ”Discussion 14” of [21] again. In other words, we can check that around the set $S(f_{0})$, $f_{0}$ is represented as the product map of a natural function and the identuty map on $S(f_{0})$ as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
We can also check Main Theorem 3 (6) easily thanks to our arguments. A result close to this is also presented in Theorem 1 (3) of [18]. There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there.
This completes the proof. ∎
Example 1.
Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$. For this we can choose $a$ as an arbitrary positive real number. The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle: ${\Delta}^{k}$ is $D^{k}$ in this case. Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3. This is also for Main Theorem 2. Main Theorem 2 gives one of most important specific cases.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example. We also apply related arguments implicitly in some of our previous arguments in the present paper.
See also Remark 4, in the next section, for example.
We give additional explicit cases as Examples. They are also a kind of remarks.
Example 2.
Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ satisfying the following.
  • The function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is a Morse function.
  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on a smooth closed manifold. The manifold is a non-singular real algebraic manifold of dimension $m-n$.
  • The minimum of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is 00 and the condition ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$ holds.
Example 1 is generalized here.
We have a special generic map $f_{0}$ as in the case of Main Theorems 1 and 2.
Example 3.
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example. See also Remark 3 of [15]. Some examples of Example 2 are regarded as specific cases of the present cases.
3. Meanings of our main result and remarks.
The first remark compares our main result to previous results by the author.
Remark 1.
In [15, 16, 18], we consider cases where polynomials are given beforehand. In short, only STEP 2 in our proof of Main Theorems 1 and 2 is considered and related short exposition on STEP 3 in our proof of Main Theorems 1 and 2 is presented where the latter presentation is a kind of short remark. This does not affect main ingredients of [15, 16, 18]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
The following is on another problem.
Remark 2.
As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing canonical projections. Some are closely related to Remark 1.
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
[41] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006, [45] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [24, 34] show explicit answers. [14] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[15] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [4, 49, 50] are also closely related, help the author to do such studies, and motivate us to study further. They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
Remark 3 is a remark on STEP 1 in our proof of Main Theorems 1 and 2. This is also independent of our present study. Here we also add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33]. See also [42]. Remark 3 is also closely related to [4]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory. This mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not present elementary notions on graphs precisely. We except related knowledge. We also note that we omit several rigorous discussions in Remark 3 and that for our main ingredients of the present paper we do not need to understand this rigorously. However, we give comments that this is also related to Remark 1 and 2.
Remark 3.
[4] considers a finite graph $K$ enjoying the following properties.
  • We can embed $K$ by a piecewise smooth embedding $e:K\rightarrow{\mathbb{R}}^{2}$.
  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is a smooth embedding.
  • At distinct vertices, the values of ${\pi}_{2,1}\circ e$ are always distinct.
  • The degree of each vertex is $1$ or $3$.
  • At the vertex where the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ has a local extremum, the degree of the vertex is $1$.
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation is considered. Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
  1. (1)
    $e(K)\subset R_{K}$.
  2. (2)
    The Poincaré-Reeb graph of $R_{K}$ is isomorphic to and naturally regarded as $e(K)$.
  3. (3)
    We also have the situation in the following. Of course $\overline{R_{K}}$ is the closure of $R_{K}$ in ${\mathbb{R}}^{2}$.
    1. (a)
      The restriction of ${\pi}_{2,1}$ or ${\pi}_{2,1}{\mid}_{\overline{R_{K}}}:\overline{R_{K}}\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
    2. (b)
      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
A desired region $R_{K}$ can be also obtained by an approximation of the Weierstrass type in the class $C^{2}$. According to arguments similar to ones in ”Special Case 5” of [21], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
The following is a remark on Main Theorem 3.
Remark 4.
We can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
The following is closely related to one of central problems in real algebraic geometry. This may be closely related to our future work.
Remark 5.
Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces has been difficult and important in real algebraic geometry. This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools. At present, even explicit construction and classifications for cases of real algebraic curves are difficult. Our pioneering studies of the construction respect unit spheres and canonical projections of them where we do not restrict attention to $1$-dimensional cases. Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.
The author recognizes that our construction shows explicit cases of Viro’s methods. In short, Viro’s methods glue local hypersurfaces together. We can have non-singular real algebraic hypersurfaces by deforming and smoothing (Viro’s patchworking). In our paper we do not need to understand this theory well. However, to study more, we need to understand the theory more.
We close our paper by a short remark on [18].
Remark 6.
[18] has no essential errors and we have also found that we need several more rigorous exposition. Some of these important arguments are discussed in the present paper. We will revise the version and submit our revised version to the refereed journal, based on the presented comments by editors and referees on [18] and our new arguments and idea in the present paper. Note also that the content of [18] is independent of the content of the present paper essentially. We have also found we can revise [18] in a self-contained way although we need to respect some of our paper. We will do so as we can.
4. Acknowledgement.
This new study is based on studies on special generic maps and explicit construction of real algebraic functions and maps of the author. One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry. See also the previous versions [19].